# Merge Kernel for Bayesian Optimization on Permutation Space

### v2

@@ Abstract
Bayesian Optimization (BO) algorithm is a standard tool for black-box optimization problems. The current state-of-the-art BO approach for permutation spaces relies on the Mallows kernel-an $\Omega(n^{2})$ representation that explicitly enumerates every pairwise comparison. Inspired by the close relationship between the Mallows kernel and pairwise comparison, we propose a novel framework for generating kernel functions on permutation space based on sorting algorithms. Within this framework, the Mallows kernel can be viewed as a special instance derived from enumeration sort. Further, we introduce the Merge Kernel constructed from merge sort, which replaces the quadratic complexity with $\Theta(n\log n)$ to achieve the lowest possible complexity. The resulting feature vector is significantly shorter, can be computed in linearithmic time, yet still efficiently captures meaningful permutation distances. To boost robustness and right-invariance without sacrificing compactness, we further incorporate three lightweight, task-agnostic descriptors: (1) a shift histogram, which aggregates absolute element displacements and supplies a global misplacement signal; (2) a split-pair line, which encodes selected long-range comparisons by aligning elements across the two halves of the whole permutation; and (3) sliding-window motifs, which summarize local order patterns that influence near-neighbor objectives. Our empirical evaluation demonstrates that the proposed kernel consistently outperforms the state-of-the-art Mallows kernel across various permutation optimization benchmarks. Results confirm that the Merge Kernel provides a more compact yet more effective solution for Bayesian optimization in permutation space.
Introduction
@@
As one of the most widely adopted approaches to black-box optimization, Bayesian optimization (BO) (Shahriari et al. 2015) has found broad application in machine-learning hyper-parameter tuning (Wu et al. 2019), financial portfolio optimization (Gonzalvez et al. 2019), chemical‐material discovery (Luo et al. 2025), and catalyst formulation design (Xie et al. 2023). BO employs probabilistic surrogate models—most commonly Gaussian processes (GPs)—to approximate the unknown objective and uses an acquisition function to balance exploration and exploitation, thereby approaching the global optimum with the fewest possible evaluations.
@@
Most research on BO to date has concentrated on continuous (Greenhill et al. 2020) and categorical design spaces (Garrido-Merchán and Hernández-Lobato 2020; Nguyen et al. 2020; Bartoli et al. 2025), whereas applications to permutation spaces remain comparatively underexplored. This gap is striking given the ubiquity of permutation optimization in both theory and practice: canonical examples include the travelling-salesperson problem (TSP), the sequencing of operations in automated experimental pipelines (Guidi, Seeberger, and Gilmore 2020), and the sequential order-of-addition experiments (Lin and Rios 2025). In addition, permutation optimization frequently arises in diverse AI applications, including scheduling tasks in robotic planning (Alatartsev, Stellmacher, and Ortmeier 2015), optimizing experimental sequences (Blau et al. 2022), and other sequential decision-making problems (Sun and Giles 2001; Wen et al. 2023). Therefore, extending Bayesian optimization to permutation spaces holds significant theoretical and practical value for artificial intelligence as well as broader academic and industrial applications.
Successfully deploying BO in permutation spaces hinges on equipping the GP surrogate with a kernel that faithfully quantifies the similarity between two permutations. Existing approaches have evolved along two principal lines: (1) General-purpose discrete BO frameworks—most notably COMBO (Oh et al. 2019). COMBO combines a graph-Laplacian kernel with local search to accommodate heterogeneous discrete variables and thus enjoys broad applicability. However, it relies on manually crafted adjacency graphs or hash encodings, which struggle to capture the comparison sequences and cyclic shifts unique to permutations; consequently, its efficiency degrades on tasks that involve frequent high-order swaps or mixed local/global reorderings. (2) Permutation-specific kernels—chief among them the Mallows kernel in BOPS(Deshwal et al. 2022). This kernel is built on the Kendall–$\tau$ distance, representing a permutation via the inversion counts generated by enumeration sort, yielding a feature dimension of $O(n^{2})$. The dimension therefore scales quadratically with $n$, and because $2^{n^{2}}\!\gg n!$, the vast majority of features fail to correspond to any valid permutation, resulting in both statistical and computational redundancy.
@@
To further enhance robustness and expressiveness, we augment the Merge kernel with three lightweight descriptors:
  • Shift Histogram – captures invariance under cyclic shifts;
  • Split-Pair Line – emphasises the relative order of elements across partition boundaries;
  • Sliding-Window Motif – extracts local sorting patterns, increasing sensitivity to local perturbations.
This design attains the shortest possible encoding while simultaneously eliminating the high-dimensional redundancy of the Mallows kernel and the structural blind spots of COMBO, thereby providing an efficient and accurate similarity measure for Bayesian optimization over permutation spaces.
@@
  • Merge kernel. Applying the framework to merge sort yields Merge Kernel, whose $O(n\log n)$ construction matches the information-theoretic lower bound on comparison complexity. We further enrich it with three lightweight descriptors to boost predictive power at minimal cost.
  • Comprehensive evaluation. We assess the effectiveness of our kernels on diverse synthetic and real-world benchmarks, consistently outperforming the state-of-the-art Mallows kernel in both accuracy and efficiency.
@@
Background and Related Works
Permutation Optimization
Here we describe the problem formulation of permutation optimization with a fixed length $n\in\mathcal{N}$. Let $[n]=\{1,2,...,n\}$, a permutation is a function $\pi:[n]\longrightarrow[n]$ such that $\pi$ is bijective. The set of all permutations of $[n]$ is the symmetric group
@@
We are given a costly-to-evaluate, possibly noisy black-box function
$f:\;\mathcal{S}_{n}\longrightarrow\mathbb{R},$
which assigns a real-valued quality (e.g., cost, loss, reward) to every permutation $\pi$. Hence, the optimization problem can be formulated as
@@
Bayesian Optimization
Bayesian optimization (BO) (Shahriari et al. 2015) is an optimization algorithm for black-box objective functions that no closed-form expression or gradient information is available and whose evaluation is often an expensive physical or computational experiment. The algorithm first fits the observed data with a surrogate model, most commonly a Gaussian process (GP) (Williams and Rasmussen 2006), and then employs an acquisition function to select the next query point, balancing exploration of uncertain regions against exploitation of promising areas identified by the surrogate.
The kernel $K(\mathbf{x},\mathbf{x}^{\prime})$ is the central design lever in a Gaussian-process surrogate: it defines the similarity metric between inputs, thereby specifying the prior smoothness assumptions and, through GP inference, the posterior mean and uncertainty. Extending BO to any new search domain is therefore tantamount to endowing that domain with an appropriate kernel function. Whereas Euclidean spaces typically rely on Gaussian (RBF) kernels, discrete structures—and permutations in particular—require bespoke constructions that faithfully encode ordering relationships. In the next subsection we revisit the Mallows kernel on the symmetric group $\mathcal{S}_{n}$, outline its computational and statistical shortcomings, and use these observations to motivate our own kernel design.
Mallows Kernel for Permutation Space
BOPS-H (Deshwal et al. 2022) is the current state-of-the-art BO algorithm for permutation optimization, which propose to employ Mallows kernel (Jiao and Vert 2015) on the symmetric group $\mathcal{S}_{n}$ in a similar manner to the RBF kernel on the Euclidean space. The Mallows kernel $K_{mal}(\pi,\pi^{\prime})$ for the permutation pair ($\pi$, $\pi^{\prime}$) is defined as the exponential negative of the number of discordant pairs $n_{d}(\pi,\pi^{\prime})$ between $\pi$ and $\pi^{\prime}$:
@@
where $l\geq 0$ is the length-scale parameter of the Mallows kernel, and $d(\pi,\pi^{\prime})$ is the Kendall-$\tau$ distance (Kendall 1938) which counts the number of pairs of elements ordered oppositely by $\pi$ and $\pi^{\prime}$:
@@
Algorithm
Generating Kernel based on Sorting Algorithm
$\Phi_{Mal}(\pi)\;\in\;\{-1,1\}^{\binom{n}{2}}$
where each coordinate corresponds to the comparison of a pair of elements: -1 if they are in ascending order, and +1 otherwise. We then have
$\begin{split}K_{Mal}(\pi,\sigma)&=\exp\!\Bigl{(}-\frac{\|\Phi_{Mal}(\pi)-\Phi_{Mal}(\sigma)\|^{2}}{2\ell^{2}}\Bigr{)}\\ &=\;k_{\mathrm{RBF}}\bigl{(}\Phi_{Mal}(\pi),\,\Phi_{Mal}(\sigma)\bigr{)}.\end{split}$ (3)
Since the Gaussian RBF kernel $k_{\mathrm{RBF}}$ is strictly positive definite on $\mathbb{R}^{d}$ and thus satisfies Mercer’s condition (Mercer 1909), and because positive definiteness is preserved under composition with any deterministic mapping $\Phi$, it follows that $K(\pi,\sigma)$ constructed from $\Phi$ also satisfies Mercer’s condition and is therefore a valid kernel function.
However, not every sorting algorithm can induce a valid feature mapping suitable for kernel construction. This is because the mapping from permutations to feature space must yield feature vectors of fixed length; otherwise, feature vectors of differing lengths would not be compatible with the RBF kernel. Hence, only sorting algorithms that have a fixed comparison path and a constant number of comparisons across all inputs can generate valid feature mappings. In fact, most sorting algorithms commonly known to have complexity of order $O(n\log n)$—such as quicksort, heapsort, and standard merge sort—do not satisfy the requirement of constant comparison counts. Nevertheless, merge sort is an exception: although typically merge sort ceases comparisons once all elements from one subsequence have been merged, redundant comparisons can be artificially introduced during the merge procedure. This ensures that both the comparison path and the number of comparisons remain identical across different permutations, thus establishing a fixed comparison path. Consequently, we select merge sort as the feature mapping $\Phi_{Mer}(\pi)$ for constructing permutation-space kernel function $K_{Mer}$, following Equation 3:
$\begin{split}K_{Mer}(\pi,\sigma)&=\exp\!\Bigl{(}-\frac{\|\Phi_{Mer}(\pi)-\Phi_{Mer}(\sigma)\|^{2}}{2\ell^{2}}\Bigr{)}\\ &=\;k_{\mathrm{RBF}}\bigl{(}\Phi_{Mer}(\pi),\,\Phi_{Mer}(\sigma)\bigr{)}.\end{split}$ (4)
The element-pair comparison mapping $\Phi_{Mer}(\pi)$ is shown in Algorithm 1.
@@
  Let $V_{Left}=\Phi_{Mer}(\pi[:mid])$
  Let $V_{Right}=\Phi_{Mer}(\pi[mid:])$
  Let $\hat{\pi}_{l}=\text{sorted}(\pi[:mid])$
  Let $\hat{\pi}_{r}=\text{sorted}(\pi[mid:])$
  Let $V_{Merge}=[]$
  Let $i=j=0$
@@
We have established that merge sort, with a specially designed fixed-comparison procedure, is uniquely capable of constructing valid kernel functions among common sorting algorithms with a complexity of $\Omega(n\log n)$. We now demonstrate that the feature vector derived from merge sort achieves the theoretical lower bound on vector length for lossless permutation encoding. First, note that the lower bound on time complexity for any comparison-based sorting algorithm is $\Omega(n\log n)$; as this complexity directly corresponds to the number of element comparisons during sorting, it similarly sets a lower bound on the length of the feature vector. On the other hand, consider the permutation space consisting of all $n!$ possible permutations of length $n$. From an information-theoretic viewpoint, encoding all $n!$ permutations without loss using a binary feature vector composed solely of ${-1,1}$ requires a minimum vector length of $\log_{2}(n!)$. Applying Stirling’s approximation (Donald et al. 1999), we have:
$\log_{2}(n!)=n\log_{2}n-nlog_{2}e+O(\log n)=\Omega(n\log n)$
Consequently, the feature vector length of the merge-sort-based kernel (Merge Kernel) reaches this theoretical lower bound for lossless permutation encoding.
Adding Invariance to Merge Kernel
It is worth noting that, when constructing the Merge kernel via merge sort, we have not required the feature mapping $\Phi$ to possess any group invariance property, such as right-invariance. Traditionally, a permutation-distance measure should be invariant under right multiplication, meaning that applying an identical right-multiplication operation to two permutations should not alter the distance between them. However, only sorting algorithms with complexity $O(n^{2})$ can yield fully right-invariant kernel functions, since such invariance necessitates exhaustive pairwise comparisons among all $\frac{n(n-1)}{2}$ pairs of elements—an impossibility for more efficient sorting algorithms like merge sort. Consequently, although the Merge kernel achieves better computational efficiency through a more compact encoding, it sacrifices a certain degree of performance due to the loss of right-invariance.
To address this issue, we can naturally consider trading off some compactness by slightly increasing the encoding length—while preserving the overall $\Omega(n\log n)$ complexity—to enhance right-invariance and thus compensate for the associated performance loss. We introduce three additional encoding strategies, whose resulting feature vectors are concatenated to the original Merge feature mapping to further improve the kernel’s performance.
Algorithm 2 Sliding-window Motif Feature Mapping $\Phi_{Sli}(\pi)$
Input: permutation $\pi$ of length $n$; window length $w$
Output: sliding-window motif vector $\Phi_{Sli}(\pi)$

### v3

# From Sorting Algorithms to Scalable Kernels: Bayesian Optimization in High-Dimensional Permutation Spaces

@@ Abstract
Bayesian Optimization (BO) is a powerful tool for black-box optimization, but its application to high-dimensional permutation spaces is severely limited by the challenge of defining scalable representations. The current state-of-the-art BO approach for permutation spaces relies on an exhaustive $\Omega(n^{2})$ pairwise comparison, inducing a dense representation that is impractical for large-scale permutations. To break this barrier, we introduce a novel framework for generating efficient permutation representations via kernel functions derived from sorting algorithms. Within this framework, the Mallows kernel can be viewed as a special instance derived from enumeration sort. Further, we introduce the Merge Kernel , which leverages the divide-and-conquer structure of merge sort to produce a compact, $\Theta(n\log n)$ to achieve the lowest possible complexity with no information loss and effectively capture permutation structure. Our central thesis is that the Merge Kernel performs competitively with the Mallows kernel in low-dimensional settings, but significantly outperforms it in both optimization performance and computational efficiency as the dimension $n$ grows. Extensive evaluations on various permutation optimization benchmarks confirm our hypothesis, demonstrating that the Merge Kernel provides a scalable and more effective solution for Bayesian optimization in high-dimensional permutation spaces, thereby unlocking the potential for tackling previously intractable problems such as large-scale feature ordering and combinatorial neural architecture search.
1 Introduction
@@
As one of the most widely adopted approaches to black-box optimization, Bayesian optimization (BO) Shahriari et al. (2015) has found broad application in machine-learning hyper-parameter tuning Wu et al. (2019), financial portfolio optimization Gonzalvez et al. (2019), chemical and material discovery Luo et al. (2025), and catalyst formulation design Xie et al. (2023). BO employs probabilistic surrogate models—most commonly Gaussian processes (GPs)—to approximate the unknown objective and uses an acquisition function to balance exploration and exploitation, thereby approaching the global optimum with the fewest possible evaluations.
@@
Most research on BO to date has concentrated on continuous Greenhill et al. (2020) and categorical design spaces Garrido-Merchán & Hernández-Lobato (2020); Nguyen et al. (2020); Bartoli et al. (2025), whereas applications to permutation spaces remain comparatively underexplored. This gap is striking given the ubiquity of permutation optimization in both theory and practice: canonical examples include the traveling-salesperson problem (TSP), the sequencing of operations in automated experimental pipelines Guidi et al. (2020), and the sequential order-of-addition experiments Lin & Rios (2025). In addition, permutation optimization frequently arises in diverse AI applications, including scheduling tasks in robotic planning Alatartsev et al. (2015), optimizing experimental sequences Blau et al. (2022), and other sequential decision-making problems Sun & Giles (2001); Wen et al. (2023). Therefore, extending Bayesian optimization to permutation spaces holds significant theoretical and practical value for artificial intelligence as well as broader academic and industrial applications.
[…]
@@
  • Merge kernel. Applying the framework to merge sort yields Merge Kernel, whose $O(n\log n)$ construction matches the information-theoretic lower bound on comparison complexity.
  • Comprehensive evaluation. We assess the effectiveness of our kernels on diverse synthetic and real-world benchmarks. Results on low-dimensional benchmarks show competitive performance against the state-of-the-art Mallows kernel, while it significantly outperforms the Mallows kernel on high-dimensional benchmarks.
@@
2 Background and Related Works
2.1 Permutation Optimization
Here we describe the problem formulation of permutation optimization with a fixed length $n\in\mathbb{N}$. Let $n]=\{1,2,...,n\}$, a permutation is a function $\pi:[n]\longrightarrow[n]$ such that $\pi$ is bijective. The set of all permutations of $[n]$ is the symmetric group
@@ 2.1 Permutation Optimization
We are given a costly-to-evaluate, possibly noisy black-box function $f:\;\mathcal{S}_{n}\longrightarrow\mathbb{R}$, which assigns a real-valued quality (e.g., cost, loss, reward) to every permutation $\pi$. Hence, the optimization problem can be formulated as
@@
2.2 Bayesian Optimization
Bayesian optimization (BO) Shahriari et al. (2015) is an optimization algorithm for black-box objective functions that no closed-form expression or gradient information is available and whose evaluation is often an expensive physical or computational experiment. The algorithm first fits the observed data with a surrogate model, most commonly a Gaussian process (GP) Williams & Rasmussen (2006), and then employs an acquisition function to select the next query point, balancing exploration of uncertain regions against exploitation of promising areas identified by the surrogate.
The kernel $K(\mathbf{x},\mathbf{x}^{\prime})$ is the central design lever in a Gaussian-process surrogate: it defines the similarity metric between inputs, thereby specifying the prior smoothness assumptions and, through GP inference, the posterior mean and uncertainty. Extending BO to any new search domain is therefore tantamount to endowing that domain with an appropriate kernel function. Whereas Euclidean spaces typically rely on Gaussian (RBF) kernels, discrete structures—and permutations in particular—require bespoke constructions that faithfully encode ordering relationships. For example, Oh et al. (2022) uses the Position kernel Zaefferer et al. (2014) and L-ensemble with Acquisition Weights to extend BO on permutation space to a batched BO scheme.
While significant progress has been made in scaling BO to high-dimensional continuous domains (e.g., Eriksson et al. (2019)Wang et al. (2016)), extending BO to large-scale structured discrete domains like permutations presents a distinct set of challenges centered on kernel design, which is the primary focus of this work.
2.3 Mallows Kernel for Permutation Space
BOPS-H Deshwal et al. (2022) is the current state-of-the-art BO algorithm for permutation optimization, which proposes to employ Mallows kernel Jiao & Vert (2015) on the symmetric group $\mathcal{S}_{n}$ in a similar manner to the RBF kernel on the Euclidean space. The Mallows kernel $K_{mal}(\pi,\pi^{\prime})$ for the permutation pair ($\pi$, $\pi^{\prime}$) is defined as the exponential negative of the number of discordant pairs $n_{d}(\pi,\pi^{\prime})$ between $\pi$ and $\pi^{\prime}$:
@@ 2.3 Mallows Kernel for Permutation Space
where $l\geq 0$ is the length-scale parameter of the Mallows kernel, and $d(\pi,\pi^{\prime})$ is the Kendall-$\tau$ distance Kendall (1938) which counts the number of pairs of elements ordered oppositely by $\pi$ and $\pi^{\prime}$:
@@
3 Merge Kernel: Generating Kernels from sorting algorithms
@@
 Let $V_{Left},V_{Right}=\Phi_{Mer}(\pi[:mid]),\Phi_{Mer}(\pi[mid:])$
 Let $\hat{\pi}_{l},\hat{\pi}_{r}=\text{sorted}(\pi[:mid]),\text{sorted}(\pi[mid:])$
 Let $V_{Merge}=[]$, $i=j=0$
@@
$\Phi_{Mal}(\pi)\;\in\;\{0,1\}^{\binom{n}{2}}$
where each coordinate corresponds to the comparison of a pair of elements: 0 if they are in ascending order, and 1 otherwise. We then have
$\begin{split}K_{Mal}(\pi,\pi^{\prime})&=\exp\!\Bigl(-\frac{\|\Phi_{Mal}(\pi)-\Phi_{Mal}(\pi^{\prime})\|^{2}}{2\ell^{2}}\Bigr)\\ &=\;K_{\mathrm{RBF}}\bigl(\Phi_{Mal}(\pi),\,\Phi_{Mal}(\pi^{\prime})\bigr).\end{split}$ (3)
under the reparameterization $l=\frac{1}{2\ell^{2}}$ for the length-scale parameter $\ell$ in the Gaussian RBF kernel. Since the RBF kernel $K_{\mathrm{RBF}}$ is strictly positive definite on $\mathbb{R}^{d}$ and thus satisfies Mercer’s condition Mercer (1909), and because positive definiteness is preserved under composition with any deterministic mapping $\Phi$, it follows that $K(\pi,\pi^{\prime})$ constructed from $\Phi$ also satisfies Mercer’s condition and is therefore a valid kernel function.
However, not every sorting algorithm can induce a valid feature mapping suitable for kernel construction. This is because the mapping from permutations to feature space must yield feature vectors of fixed length; otherwise, feature vectors of differing lengths would not be compatible with the RBF kernel. Hence, only sorting algorithms that have a fixed comparison path and a constant number of comparisons across all inputs can generate valid feature mappings, e.g., a fixed sorting network with a predetermined comparator sequence Batcher (1968). In fact, most sorting algorithms commonly known to have complexity of order $O(n\log n)$—such as quicksort, heapsort, and standard merge sort—do not satisfy the requirement of constant comparison counts. Nevertheless, merge sort could be an exception: although typically merge sort ceases comparisons once all elements from one subsequence have been merged, redundant comparisons can be artificially introduced during the merge procedure, resulting in a fixed number of comparisons L+R1L+R-1, where $L$ and $R$ are the length of the two subsequences. This ensures that both the comparison path and the number of comparisons remain identical across different permutations, thus establishing a fixed comparison path. Consequently, we select merge sort as the feature mapping $\Phi_{Mer}(\pi)$ for constructing permutation-space kernel function $K_{Mer}$, following Equation 3:
$\begin{split}K_{Mer}(\pi,\pi^{\prime})&=\exp\!\Bigl(-\frac{\|\Phi_{Mer}(\pi)-\Phi_{Mer}(\pi^{\prime})\|^{2}}{2\ell^{2}}\Bigr)\\ &=\;K_{\mathrm{RBF}}\bigl(\Phi_{Mer}(\pi),\,\Phi_{Mer}(\pi^{\prime})\bigr).\end{split}$ (4)
The element-pair comparison mapping $\Phi_{Mer}(\pi)$ is shown in Algorithm 1. Here we present an example for permutation [1,4,3,2] to show the Merging process.
Example: Feature Mapping for $\pi=(1,4,3,2)$
  1. 1.
    Initial Split: $\pi$ splits into $L_{0}=(1,4)$ and $R_{0}=(3,2)$.
  2. 2.
    Recurse on $L_{0}=(1,4)$: Merging sorted lists ‘[1]’ and ‘[4]’ yields feature vector $V_{L}=[0]$.
  3. 3.
    Recurse on $R_{0}=(3,2)$: Merging sorted lists ‘[2]’ and ‘[3]’ yields feature vector $V_{R}=[1]$.
  4. 4.
    Final Merge: Merge sorted lists ‘(1,4)’ and ‘(2,3)’. The fixed comparison path generates the merge vector $V_{Merge}=[0(1<2),1(4>2),1(4>3),1(\text{left padding})]$.
  5. 5.
    Concatenate: The final feature vector is $\Phi_{Mer}(\pi)=V_{L}\oplus V_{R}\oplus V_{Merge}=[0]\oplus[1]\oplus[0,1,1,1]=[0,1,0,1,1,1]$.
@@
We have established that merge sort, with a specially designed fixed-comparison procedure, is uniquely capable of constructing valid kernel functions among common sorting algorithms with a complexity of $\Omega(n\log n)$. We now demonstrate that the feature vector derived from merge sort achieves the theoretical lower bound on vector length for lossless permutation encoding. First, note that the lower bound on time complexity for any comparison-based sorting algorithm is $\Omega(n\log n)$; as this complexity directly corresponds to the number of element comparisons during sorting, it similarly sets a lower bound on the length of the feature vector. On the other hand, consider the permutation space consisting of all $n!$ possible permutations of length $n$. From an information-theoretic viewpoint, encoding all $n!$ permutations without loss using a binary feature vector composed solely of 0,1{0,1} requires a minimum vector length of $\log_{2}(n!)$. Applying Stirling’s approximation Donald et al. (1999), we have: $\log_{2}(n!)=n\log_{2}n-nlog_{2}e+O(\log n)=\Omega(n\log n)$. Consequently, the feature vector length of the merge-sort-based kernel (Merge Kernel) reaches this theoretical lower bound for lossless permutation encoding.